# On the Geometric Coherence of Global Aggregation in Federated GNN

### v1

@@ Abstract
Federated Learning (FL) enables distributed training across multiple clients without centralized data sharing, while Graph Neural Networks (GNNs) model relational data through message passing. In federated GNN settings, client graphs often exhibit heterogeneous structural and propagation characteristics. When standard aggregation mechanisms are applied to such heterogeneous updates, the global model may converge numerically while exhibiting degraded relational behavior.Our work identifies a geometric failure mode of global aggregation in Cross- Domain Federated GNNs. Although GNN parameters are numerically represented as vectors, they encode relational transformations that govern the direction, strength, and sensitivity of information flow across graph neighborhoods. Aggregating updates originating from incompatible propagation regimes can therefore introduce destructive interference in this transformation space.This leads to loss of coherence in global message passing. Importantly, this degradation is not necessarily reflected in conventional metrics such as loss or accuracy.To address this issue, we propose GGRS (Global Geometric Reference Structure), a server-side framework that regulates client updates prior to aggregation based on geometric admissibility criteria. GGRS preserves directional consistency of relational transformations as well as maintains diversity of admissible propagation subspaces. It also stabilizes sensitivity to neighborhood interactions, without accessing client data or graph topology. Experiments on heterogeneous GNN-native, Amazon Co-purchase datasets demonstrate that GGRS preserves global message-passing coherence across training rounds by highlighting the necessity of geometry-aware regulation in federated graph learning.
@@
Federated Learning (FL) addresses distributed optimization problems in which a shared model is trained across multiple decentralized clients under communication and data-access constraints [1]. At each communication round, locally optimized parameter updates are aggregated at a global server to form a new global model, which is subsequently broadcast to clients for further local optimization.Whearas, whearas Cross-domain federated machine learning [4] refers to a federated setting in which participating clients hold data originating from distinct domains or distributions, rather than being non-IID partitions of a single dataset. In this regime, each client’s local data distribution may differ substantially in feature statistics, label semantics, or structural characteristics, leading to stronger heterogeneity than conventional FL. The goal remains to learn a shared global model via decentralized optimization, but aggregation becomes more challenging due to domain-induced misalignment across client updates. Classical aggregation schemes such as Federated Averaging (FedAvg) implicitly assume that client updates lie in a common parameter space where linear combination preserves model semantics.GNN[5] fundamentally violate this assumption. A GNN does not merely define a parameterized predictor, but rather a parameterized message-passing operator acting on graph-structured domains [5, 6]. For a fixed architecture, learnable parameters induce a family of linear and nonlinear operators that govern how node representations are updated as a function of neighborhood structure, aggregation depth, and interaction strength. Consequently, parameter updates in GNNs correspond to perturbations of an operator that determines the geometry of information flow over the graph, rather than independent numerical adjustments.
@@
So in the further process, the global model produced by the server defines the initialization for all clients in the subsequent round. If this model encodes a distorted message-passing operator, local optimization must first compensate for this distortion before adapting to the local graph structure. This induces corrective gradients that are misaligned across clients, increasing update variance and amplifying destructive interference at the next aggregation step. Over multiple rounds, this feedback loop leads to progressive destabilization of the global operator, even as standard convergence criteria are satisfied.Existing approaches to federated GNN learning primarily address heterogeneity through personalization, client clustering, or adaptive weighting strategies[10, 2]. While effective in mitigating performance degradation, these methods do not explicitly characterize aggregation as an operation on message-passing operators, nor do they impose constraints that preserve the geometric integrity of the induced transformations. As a result, failure modes arising from operator-level incompatibility rather than optimization error remain unaddressed.
@@ 1.1  Geometric Observations in Federated GNN
From our study, we have summarized a set of fundamental observations regarding the geometric behavior of GNN under federated aggregation. These observations are not specific to a particular architecture or optimization strategy, but arise from the interaction between graph-dependent message passing and parameter aggregation across heterogeneous clients.Our first observation is Parameters as Structured Transformations: Although model parameters are numerically represented as vectors or matrices, learning behavior does not arise from individual scalar values but from coordinated changes across parameter groups. In GNNs, parameters jointly define structured transformations that govern how information propagates across graph neighborhoods. Consequently, the semantic meaning of a parameter update is determined by its direction and interaction with other parameters, rather than by its magnitude alone. In our next inference, we saw that GNNs Learn Message-Passing Operators: A GNN induces a parameterized message-passing operator whose repeated application updates node representations as a function of graph connectivity. In a fixed architecture, parameters specify how neighborhood information is aggregated, how far information propagates, and how sensitive node embeddings are to structural perturbations. Parameter updates therefore correspond to perturbations of this operator rather than independent numerical adjustments.Next we see,Graph-Dependent Operator Geometry: We know that the operator induced by a GNN is graph-dependent. The same parameter configuration can induce distinct transformations when applied to graphs with different adjacency structures or feature distributions. As a result, the Jacobian of the GNN with respect to its parameters depends on the underlying graph, implying that parameter updates computed on different clients generally inhabit distinct operator tangent spaces.Across the literature survey, we observed some implicit Assumptions of Federated Averaging. In Standard federated aggregation mechanisms implicitly assume that client updates are commensurate in a shared Euclidean parameter space, such that linear combination preserves model semantics. This assumption is violated in federated GNNs, where client updates encode perturbations of graph-dependent operators. Linear aggregation therefore mixes updates without enforcing compatibility of their induced transformations. Finally, we assessed the Numerical Convergence Versus Geometric Coherence. Federated training can exhibit numerical convergence in terms of loss reduction or gradient norms while simultaneously degrading the geometric coherence of the message-passing operator. Such degradation may manifest as diminished sensitivity to neighborhood structure, collapse of effective propagation depth, or contraction of representations into low-dimensional subspaces.
@@ 1.2  Limitations of Accuracy-Centric Evaluation in Federated GNNs
Machine learning systems are commonly evaluated using task-level metrics such as accuracy, F1-score, or loss. These metrics implicitly assume that improvements or degradation in model performance directly reflect the quality of the learned representations. While this assumption holds reasonably well for centralized learning on i.i.d. data, it becomes insufficient in federated and graph-based learning settings, where model behavior depends critically on relational structure and iterative aggregation dynamics.Primary insight is into accuracy as a Performance Proxy in Conventional Learning Models. In conventional supervised learning, accuracy measures the alignment between model predictions and ground-truth labels on a fixed data distribution. Optimization procedures are designed to minimize empirical risk, and convergence in loss is typically interpreted as convergence in representation quality. Under these assumptions, parameter averaging or gradient-based updates preserve semantic meaning, and degradation in performance is readily observable through scalar metrics.There are some notable limitations of Accuracy Metrics in GNNs.GNNs differ fundamentally from standard models in that predictions are produced through iterative message passing over graph structure. The expressive power of a GNN depends not only on parameter values, but on how those parameters govern information propagation across neighborhoods, graph depth, and connectivity patterns. As a result, a GNN can exhibit stable accuracy while its internal message-passing dynamics degrade.In particular, accuracy does not capture whether:neighborhood information is meaningfully integrated or collapsed,long-range dependencies are preserved or suppressed,node representations remain sensitive to structural perturbations,propagation depth remains effective across layers.These properties are geometric and operator-level in nature, and their degradation may remain invisible to scalar evaluation metrics.Another inference is on accuracy Degradation in Federated Settings.In FL, accuracy is further confounded by the absence of a centralized validation distribution. Each client evaluates performance on its local data, and the global model is rarely assessed on a distribution that reflects the union of client graph structures. As a result, a global model may appear to perform adequately on average while being misaligned with all individual client distributions. Moreover, federated training introduces an iterative feedback loop: the global model is repeatedly broadcast to clients and used as the initialization for local optimization. Even small distortions in model behavior introduced by aggregation can compound over rounds, leading to progressive degradation that is not immediately reflected in early-round accuracy measurements.Empirically, federated GNNs under heterogeneous graph distributions exhibit several failure modes that are poorly explained by accuracy alone:Over-smoothing acceleration: node representations collapse into low-rank subspaces more rapidly than in centralized training.Over-squashing amplification: long-range information is increasingly compressed, reducing the effective receptive field.Update oscillation: local updates increasingly counteract global updates across rounds.Representation drift: the semantic meaning of intermediate embeddings changes across rounds despite stable task loss.These effects indicate degradation of the learned message-passing operator rather than simple underfitting or overfitting.Root Cause can be traced as Aggregation-Induced Geometric Distortion i.e. the common factor underlying these failure modes is the interaction of heterogeneous client updates at the global server.Client updates encode graph-dependent operator perturbations. Standard aggregation mechanisms combine these updates in parameter space without enforcing compatibility of their induced transformations.This process can converge numerically while simultaneously distorting the geometry of the global message-passing operator. Such distortion alters propagation directionality, collapses admissible subspaces, and destabilizes sensitivity to graph structure—effects that directly impair relational learning but may not immediately impact prediction accuracy. These observations indicate that federated GNN training cannot be reliably assessed or stabilized using accuracy-centric metrics alone. Scalar performance measures fail to diagnose degradation of message-passing geometry and provide no insight into the stability of relational representations across communication rounds.Consequently, effective federated GNN systems require mechanisms that explicitly preserve operator-level geometric coherence during aggregation. Without such mechanisms, improvements in optimization or personalization may address symptoms while leaving the underlying structural failure unresolved.
1.3  Problem Formulation and Geometric Failure in Federated GNN Aggregation
1.3.1 Federated GNN Learning Setup
@@
We consider a FL system consisting of $K$ clients and a central server. Each client $k\in\{1,\dots,K\}$ observes a local graph
$\mathcal{G}_{k}=(\mathcal{V}_{k},\mathcal{E}_{k},\mathbf{X}_{k},\mathbf{Y}_{k}),$ (1)
where $\mathcal{V}_{k}$ and $\mathcal{E}_{k}$ denote the node and edge sets, $\mathbf{X}_{k}\in\mathbb{R}^{|\mathcal{V}_{k}|\times d}$ the node features, and $\mathbf{Y}_{k}$ the node labels. Graphs are assumed to be drawn from client-specific distributions $\mathcal{G}_{k}\sim\mathcal{D}_{k}$, with $\mathcal{D}_{k}\neq\mathcal{D}_{k^{\prime}}$ in general.
@@
All clients share a common GNN architecture parameterized by $\theta\in\mathbb{R}^{p}$. Local training at client $k$ minimizes an empirical risk
$\theta_{k}^{(t)}=\arg\min_{\theta}\;\mathbb{E}_{\mathcal{G}_{k}\sim\mathcal{D}_{k}}\left[\mathcal{L}\left(f_{\theta}(\mathcal{G}_{k}),\mathbf{Y}_{k}\right)\right],$ (2)
where $f_{\theta}$ denotes the GNN predictor and $t$ indexes the communication round.
@@
After local optimization, clients transmit parameter updates $\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)}$ to the server, which computes a global update via aggregation, typically
$\theta^{(t+1)}=\theta^{(t)}+\sum_{k=1}^{K}w_{k}\Delta\theta_{k}^{(t)},$ (3)
where $w_{k}$ are aggregation weights.
1.3.2 GNNs as Parameterized Message-Passing Operators
A $L$-layer GNN can be written as the composition of message-passing operators
$\mathbf{H}^{(l+1)}=\Phi_{\theta^{(l)}}\left(\mathbf{H}^{(l)},\mathcal{G}\right),\quad l=0,\dots,L-1,$ (4)
where $\mathbf{H}^{(0)}=\mathbf{X}$ and $\Phi_{\theta^{(l)}}$ denotes a neighborhood aggregation operator. For instance, in graph convolutional networks,
$\Phi_{\theta}(\mathbf{H},\mathcal{G})=\sigma\left(\tilde{\mathbf{A}}\mathbf{H}\mathbf{W}_{\theta}\right),$ (5)
with normalized adjacency $\tilde{\mathbf{A}}$ and learnable weights $\mathbf{W}_{\theta}$.
Crucially, $\theta$ does not merely parameterize a function mapping features to labels, but induces a linear operator on node representations whose spectral and geometric properties determine information propagation across the graph.
1.3.3 Geometric Representation of Message Passing
Let $\mathcal{T}_{\theta}^{\mathcal{G}}$ denote the message-passing operator induced by parameters $\theta$ on graph $\mathcal{G}$. Linearizing around $\theta$ yields
$\mathcal{T}_{\theta+\delta}^{\mathcal{G}}\approx\mathcal{T}_{\theta}^{\mathcal{G}}+\mathcal{J}_{\theta}^{\mathcal{G}}\delta,$ (6)
where $\mathcal{J}_{\theta}^{\mathcal{G}}$ is the Jacobian of the operator with respect to parameters.
Each client update $\Delta\theta_{k}$ therefore induces a perturbation
$\Delta\mathcal{T}_{k}\;=\;\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k},$ (7)
which lies in an operator space whose geometry depends on $\mathcal{G}_{k}$.
We emphasize that $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}$ differs across clients due to differences in adjacency structure and feature statistics. Consequently, even if $\Delta\theta_{k}$ are comparable in parameter space, their induced operator perturbations $\Delta\mathcal{T}_{k}$ may be geometrically incompatible.
1.3.4 Geometric Incompatibility Under Aggregation
Standard aggregation computes
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\Delta\theta_{k},$ (8)
which induces a global operator perturbation
$\Delta\mathcal{T}^{\text{global}}=\sum_{k=1}^{K}w_{k}\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k}.$ (9)
Hence,unless the Jacobians $\{\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\}$ share aligned dominant eigenspaces or consistent sensitivity profiles, the resulting perturbation does not correspond to a coherent update of any admissible message-passing operator. Instead, opposing operator directions may cancel or distort one another, leading to a contraction or degeneration of the effective transformation.Geometrically, this corresponds to projecting updates from distinct tangent spaces onto a shared parameter space without enforcing compatibility of their induced operator geometry.We discuss certain consequences for Prediction and Iterative Training which infers,in a GNN relies on repeated application of $\mathcal{T}_{\theta}^{\mathcal{G}}$. When aggregation produces a $\theta^{(t+1)}$ whose induced operator lacks coherent spectral or directional structure, node representations converge toward low-rank or insensitive subspaces. This leads to diminished dependence of predictions on graph structure.Moreover, the federated training process forms a feedback loop. The geometrically distorted $\theta^{(t+1)}$ is broadcast to clients and serves as initialization for the next local optimization phase. Each client must then apply corrective updates to restore operator geometry suitable for its own graph distribution. These corrective updates are again geometrically misaligned across clients, amplifying interference at the next aggregation step and accelerating global operator degradation.Hence we arrive to our problem statement,i.e.,given a FLsystem with heterogeneous graph distributions $\{\mathcal{D}_{k}\}$, standard aggregation fails to preserve the geometric coherence of the induced message-passing operator. In such a case, how global aggregation can be performed such that the induced message-passing operator remains geometrically coherent—preserving directional alignment, subspace structure, and sensitivity across heterogeneous client graphs.
    ormal Derivation of Geometric Mismatch in Federated GNN Aggregation
Figure 2: Horizontal schematic of the GGRS server-side geometric aggregation mechanism.
2.1  Federated Aggregation as Parameter-Space Averaging
Let $\theta\in\mathbb{R}^{p}$ denote the parameters of a fixed GNN architecture. In standard FL, local optimization at client $k$ produces an update
$\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)},$ (10)
where $\theta^{(t)}$ is the global model at communication round $t$. The server computes the next global model via linear aggregation:
$\theta^{(t+1)}=\theta^{(t)}+\sum_{k=1}^{K}w_{k}\Delta\theta_{k}^{(t)},$ (11)
with aggregation weights $w_{k}\geq 0$, $\sum_{k}w_{k}=1$.
Equation (11) implicitly assumes that the parameter space $\mathbb{R}^{p}$ constitutes a shared, semantically meaningful vector space in which linear combination preserves model behavior.
2.2  GNN Parameters Induce Graph-Dependent Operators
Consider a GNN layer expressed as a message-passing operator
$\mathbf{H}^{(l+1)}=\Phi_{\theta^{(l)}}(\mathbf{H}^{(l)},\mathcal{G}),$ (12)
where $\mathcal{G}$ denotes the underlying graph. Let $\mathcal{T}_{\theta}^{\mathcal{G}}$ denote the composite operator induced by stacking all layers.
Crucially, $\mathcal{T}_{\theta}^{\mathcal{G}}$ depends on both $\theta$ and $\mathcal{G}$. For two distinct graphs $\mathcal{G}_{a}\neq\mathcal{G}_{b}$, the same parameter vector $\theta$ induces distinct operators
$\mathcal{T}_{\theta}^{\mathcal{G}_{a}}\neq\mathcal{T}_{\theta}^{\mathcal{G}_{b}}.$ (13)
Thus, parameters do not correspond to a single global operator, but to a family of graph-dependent operators.
2.3  Local Updates as Operator Perturbations
Let $\theta^{(t)}$ be the global model at round $t$. For client $k$, a local update $\Delta\theta_{k}$ induces a perturbation of the message-passing operator:
$\mathcal{T}_{\theta^{(t)}+\Delta\theta_{k}}^{\mathcal{G}_{k}}\approx\mathcal{T}_{\theta^{(t)}}^{\mathcal{G}_{k}}+\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\Delta\theta_{k},$ (14)
where $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}$ is the Jacobian of the operator with respect to $\theta$ evaluated on graph $\mathcal{G}_{k}$.
The term $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\Delta\theta_{k}$ represents the directional effect of the update on message passing. Since $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}$ depends on $\mathcal{G}_{k}$, operator perturbations induced by different clients lie in distinct tangent spaces.
2.4  Mismatch Induced by Linear Aggregation
Under FedAvg, the aggregated parameter update is
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\Delta\theta_{k}.$ (15)
The induced operator perturbation on graph $\mathcal{G}$ is

### v2

# On the Geometric Coherence of Global Aggregation in Federated Graph Neural Networks

@@ Abstract
Federated Learning (FL) enables distributed training across multiple clients without centralized data sharing, while Graph Neural Networks (GNNs) model relational data through message passing. In federated GNN settings, client graphs often exhibit heterogeneous structural and propagation characteristics. When standard aggregation mechanisms are applied to such heterogeneous updates, the global model may converge numerically while exhibiting degraded relational behavior. Our work identifies a geometric failure mode of global aggregation in Cross- Domain Federated GNNs. Although GNN parameters are numerically represented as vectors, they encode relational transformations that govern the direction, strength, and sensitivity of information flow across graph neighborhoods. Aggregating updates originating from incompatible propagation regimes can therefore introduce destructive interference in this transformation space. This leads to loss of coherence in global message passing. Importantly, this degradation is not necessarily reflected in conventional metrics such as loss or accuracy. To address this issue, we propose GGRS (Global Geometric Reference Structure), a server-side framework that regulates client updates prior to aggregation based on geometric admissibility criteria. GGRS preserves directional consistency of relational transformations as well as maintains diversity of admissible propagation subspaces. It also stabilizes sensitivity to neighborhood interactions, without accessing client data or graph topology. Experiments on six public benchmark graphs spanning citation, co-purchase, and co-authorship domains demonstrate that GGRS preserves global message-passing coherence across training rounds, highlighting the necessity of geometry-aware regulation in federated graph learning.
@@
Federated Learning (FL) addresses distributed optimization problems in which a shared model is trained across multiple decentralized clients under communication and data-access constraints [1]. At each communication round, locally optimized parameter updates are aggregated at a global server to form a new global model, which is subsequently broadcast to clients for further local optimization. Cross-domain federated machine learning [4] refers to a federated setting in which participating clients hold data originating from distinct domains or distributions, rather than being non-IID partitions of a single dataset. In this regime, each client’s local data distribution may differ substantially in feature statistics, label semantics, or structural characteristics, leading to stronger heterogeneity than conventional FL. The goal remains to learn a shared global model via decentralized optimization, but aggregation becomes more challenging due to domain-induced misalignment across client updates. Classical aggregation schemes such as Federated Averaging (FedAvg) implicitly assume that client updates lie in a common parameter space where linear combination preserves model semantics. GNNs [5] fundamentally violate this assumption. A GNN does not merely define a parameterized predictor, but rather a parameterized message-passing operator acting on graph-structured domains [5, 6]. For a fixed architecture, learnable parameters induce a family of linear and nonlinear operators that govern how node representations are updated as a function of neighborhood structure, aggregation depth, and interaction strength. Consequently, parameter updates in GNNs correspond to perturbations of an operator that determines the geometry of information flow over the graph, rather than independent numerical adjustments.
@@
The global model produced by the server defines the initialization for all clients in the subsequent round. If this model encodes a distorted message-passing operator, local optimization must first compensate for this distortion before adapting to the local graph structure. This induces corrective gradients that are misaligned across clients, increasing update variance and amplifying destructive interference at the next aggregation step. Over multiple rounds, this feedback loop leads to progressive destabilization of the global operator, even as standard convergence criteria are satisfied. Existing approaches to federated GNN learning primarily address heterogeneity through personalization, client clustering, or adaptive weighting strategies [2, 14, 13]. While effective in mitigating performance degradation, these methods do not explicitly characterize aggregation as an operation on message-passing operators, nor do they impose constraints that preserve the geometric integrity of the induced transformations. As a result, failure modes arising from operator-level incompatibility rather than optimization error remain unaddressed.
@@ 1.1  Geometric Observations in Federated GNN
From our study, we have summarized a set of fundamental observations regarding the geometric behavior of GNN under federated aggregation. These observations are not specific to a particular architecture or optimization strategy, but arise from the interaction between graph-dependent message passing and parameter aggregation across heterogeneous clients.
Parameters as Structured Transformations: Although model parameters are numerically represented as vectors or matrices, learning behavior does not arise from individual scalar values but from coordinated changes across parameter groups. In GNNs, parameters jointly define structured transformations that govern how information propagates across graph neighborhoods. Consequently, the semantic meaning of a parameter update is determined by its direction and interaction with other parameters, rather than by its magnitude alone.
GNNs Learn Message-Passing Operators: A GNN induces a parameterized message-passing operator whose repeated application updates node representations as a function of graph connectivity. In a fixed architecture, parameters specify how neighborhood information is aggregated, how far information propagates, and how sensitive node embeddings are to structural perturbations. Parameter updates therefore correspond to perturbations of this operator rather than independent numerical adjustments.
Graph-Dependent Operator Geometry: The operator induced by a GNN is graph-dependent. The same parameter configuration can induce distinct transformations when applied to graphs with different adjacency structures or feature distributions. As a result, the Jacobian of the GNN with respect to its parameters depends on the underlying graph, implying that parameter updates computed on different clients generally inhabit distinct operator tangent spaces.
Implicit Assumptions of Federated Averaging. Standard federated aggregation mechanisms implicitly assume that client updates are commensurate in a shared Euclidean parameter space, such that linear combination preserves model semantics. This assumption is violated in federated GNNs, where client updates encode perturbations of graph-dependent operators. Linear aggregation therefore mixes updates without enforcing compatibility of their induced transformations.
Numerical Convergence Versus Geometric Coherence. Federated training can exhibit numerical convergence in terms of loss reduction or gradient norms while simultaneously degrading the geometric coherence of the message-passing operator. Such degradation may manifest as diminished sensitivity to neighborhood structure, collapse of effective propagation depth, or contraction of representations into low-dimensional subspaces.
@@ 1.2  Limitations of Accuracy-Centric Evaluation in Federated GNNs
Machine learning systems are commonly evaluated using task-level metrics such as accuracy, F1-score, or loss. These metrics implicitly assume that improvements or degradation in model performance directly reflect the quality of the learned representations. While this assumption holds reasonably well for centralized learning on i.i.d. data, it becomes insufficient in federated and graph-based learning settings, where model behavior depends critically on relational structure and iterative aggregation dynamics.
Accuracy as a Performance Proxy in Conventional Learning. In conventional supervised learning, accuracy measures the alignment between model predictions and ground-truth labels on a fixed data distribution. Optimization procedures are designed to minimize empirical risk, and convergence in loss is typically interpreted as convergence in representation quality. Under these assumptions, parameter averaging or gradient-based updates preserve semantic meaning, and degradation in performance is readily observable through scalar metrics.
Limitations of Accuracy Metrics in GNNs. GNNs differ fundamentally from standard models in that predictions are produced through iterative message passing over graph structure. The expressive power of a GNN depends not only on parameter values, but on how those parameters govern information propagation across neighborhoods, graph depth, and connectivity patterns. As a result, a GNN can exhibit stable accuracy while its internal message-passing dynamics degrade. In particular, accuracy does not capture whether:neighborhood information is meaningfully integrated or collapsed,long-range dependencies are preserved or suppressed,node representations remain sensitive to structural perturbations,propagation depth remains effective across layers. These properties are geometric and operator-level in nature, and their degradation may remain invisible to scalar evaluation metrics.
@@
Accuracy Degradation in Federated Settings. In FL, accuracy is further confounded by the absence of a centralized validation distribution. Each client evaluates performance on its local data, and the global model is rarely assessed on a distribution that reflects the union of client graph structures. As a result, a global model may appear to perform adequately on average while being misaligned with all individual client distributions. Moreover, federated training introduces an iterative feedback loop: the global model is repeatedly broadcast to clients and used as the initialization for local optimization. Even small distortions in model behavior introduced by aggregation can compound over rounds, leading to progressive degradation that is not immediately reflected in early-round accuracy measurements. Empirically, federated GNNs under heterogeneous graph distributions exhibit several failure modes that are poorly explained by accuracy alone:Over-smoothing acceleration: node representations collapse into low-rank subspaces more rapidly than in centralized training. Over-squashing amplification: long-range information is increasingly compressed, reducing the effective receptive field. Update oscillation: local updates increasingly counteract global updates across rounds. Representation drift: the semantic meaning of intermediate embeddings changes across rounds despite stable task loss. These effects indicate degradation of the learned message-passing operator rather than simple underfitting or overfitting.
Root Cause: Aggregation-Induced Geometric Distortion. The common factor underlying these failure modes is the interaction of heterogeneous client updates at the global server. Client updates encode graph-dependent operator perturbations. Standard aggregation mechanisms combine these updates in parameter space without enforcing compatibility of their induced transformations. This process can converge numerically while simultaneously distorting the geometry of the global message-passing operator. Such distortion alters propagation directionality, collapses admissible subspaces, and destabilizes sensitivity to graph structure—effects that directly impair relational learning but may not immediately impact prediction accuracy. These observations indicate that federated GNN training cannot be reliably assessed or stabilized using accuracy-centric metrics alone. Scalar performance measures fail to diagnose degradation of message-passing geometry and provide no insight into the stability of relational representations across communication rounds. Consequently, effective federated GNN systems require mechanisms that explicitly preserve operator-level geometric coherence during aggregation. Without such mechanisms, improvements in optimization or personalization may address symptoms while leaving the underlying structural failure unresolved.
    roblem Formulation and Geometric Failure in Federated GNN Aggregation
2.1  Federated GNN Learning Setup
@@
We consider a federated learning (FL) system comprising $K$ clients and a central server. Each client $k\in[K]\triangleq\{1,\dots,K\}$ holds a private graph
$\mathcal{G}_{k}=\bigl(\mathcal{V}_{k},\,\mathcal{E}_{k},\,\mathbf{X}_{k}\in\mathbb{R}^{n_{k}\times d},\,\mathbf{Y}_{k}\bigr),$ (1)
drawn from a client-specific distribution $\mathcal{D}_{k}$. Heterogeneity is encoded by the assumption $\mathcal{D}_{k}\neq\mathcal{D}_{k^{\prime}}$ for $k\neq k^{\prime}$, arising from structural differences in graph topology, feature statistics, and label distributions.
@@
All clients share a GNN architecture with parameter vector $\theta\in\mathbb{R}^{p}$. At communication round $t$, local training at client $k$ solves
$\theta_{k}^{(t)}\;=\;\arg\min_{\theta}\;\mathbb{E}_{\mathcal{G}_{k}\sim\mathcal{D}_{k}}\bigl[\mathcal{L}(f_{\theta}(\mathcal{G}_{k}),\,\mathbf{Y}_{k})\bigr].$ (2)
@@
Clients transmit parameter updates $\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)}$ to the server, which forms the next global model via weighted aggregation:
$\theta^{(t+1)}=\theta^{(t)}+\sum_{k=1}^{K}w_{k}\Delta\theta_{k}^{(t)},\qquad w_{k}\geq 0,\quad\sum_{k}w_{k}=1.$ (3)
2.2  GNN Parameters as Graph-Dependent Operator Families
An $L$-layer GNN is realized by iterated message-passing:
$\mathbf{H}^{(l+1)}=\Phi_{\theta^{(l)}}\!\bigl(\mathbf{H}^{(l)},\mathcal{G}\bigr),\quad l=0,\dots,L-1,\quad\mathbf{H}^{(0)}=\mathbf{X}.$ (4)
For a graph convolutional layer, $\Phi_{\theta}(\mathbf{H},\mathcal{G})=\sigma(\tilde{\mathbf{A}}\mathbf{H}\mathbf{W}_{\theta})$, where $\tilde{\mathbf{A}}=\hat{\mathbf{D}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{D}}^{-1/2}$ is the symmetrically-normalized adjacency with self-loops. Let $\mathcal{T}_{\theta}^{\mathcal{G}}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n\times c}$ denote the full $L$-layer operator induced by $(\theta,\mathcal{G})$.
A key structural observation is that $\theta$ does not specify a single fixed operator but rather an entire family $\{\mathcal{T}_{\theta}^{\mathcal{G}}\}_{\mathcal{G}}$, one member per graph. For distinct client graphs $\mathcal{G}_{a}\neq\mathcal{G}_{b}$,
$\mathcal{T}_{\theta}^{\mathcal{G}_{a}}\;\neq\;\mathcal{T}_{\theta}^{\mathcal{G}_{b}},$ (5)
even though $\theta$ is identical. Consequently, the parameter space $\mathbb{R}^{p}$ is not a neutral vector space: directions in $\mathbb{R}^{p}$ carry different operational meaning depending on which graph they act upon.
2.3  Operator Perturbations and Geometric Incompatibility
For client $k$, a first-order expansion around $\theta^{(t)}$ gives
$\mathcal{T}_{\theta^{(t)}+\Delta\theta_{k}}^{\mathcal{G}_{k}}\;\approx\;\mathcal{T}_{\theta^{(t)}}^{\mathcal{G}_{k}}+\underbrace{\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\,\Delta\theta_{k}}_{\displaystyle\Delta\mathcal{T}_{k}},$ (6)
where $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\in\mathbb{R}^{(n_{k}c)\times p}$ is the Jacobian of the vectorized output with respect to $\theta$ on graph $\mathcal{G}_{k}$. The term $\Delta\mathcal{T}_{k}\triangleq\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\Delta\theta_{k}$ is the induced operator perturbation: it lives in the tangent space of the operator manifold at $\mathcal{T}_{\theta^{(t)}}^{\mathcal{G}_{k}}$.
Equation (3) aggregates parameter vectors, yielding a global operator perturbation on any evaluation graph $\mathcal{G}$: